# Towards Scalable Bayesian Optimization via Gradient-Informed Bayesian Neural Networks

### v1

@@ 1 Introduction
Bayesian optimization (BO) (Močkus, 1974) has been a powerhouse for optimization of functions with unknown or partially known structure in diverse decision-making applications (Frazier, 2018). BO tackles black-box optimization problems by sequentially solving easier optimization sub-problems. This is done by combining a probabilistic surrogate model, representing the objective as a function of decision inputs, and an acquisition function (AF) that, based on the current belief about the objective, suggests points in the design space to query the objective.
BO is traditionally a zeroth-order optimization method, exclusively relying on objective function observations to refine its belief of the objective function. As for surrogate modeling of the objective, Gaussian process (GP) regression is a popular choice (Williams and Rasmussen, 2006). Recent work has demonstrated the benefits of leveraging derivative information in BO via GPs; fully observed gradients can yield $D+1$ extra observations, where $D$ is the input dimensionality, per system query at a given input. This provides useful local information around the observed points, leading to a more accurate posterior for objective function approximation (Wu et al., 2017b). Gradients have been shown to improve regret bounds in BO (Shekhar and Javidi, 2021), or used for combined global and local search in more complex, potentially multi-objective tasks (Ip et al., 2024). A standard approach of exploiting gradients (or higher-order derivatives) in BO with GPs is by defining a joint kernel that utilizes both function and gradient observations (Eriksson et al., 2018; Wu et al., 2017a). Alternatively, to mitigate challenges due to noisy function and derivative observations, gradients can be directly modeled using a probabilistic model and utilized to enforce optimality conditions in the AF (Penubothula et al., 2021; Makrygiorgos et al., 2023a, b).
@@ 2.2 Bayesian Neural Networks
[…]
@@ 4 Case Studies
We define $\mathcal{R}_{t}$ as the instantaneous regret at iteration $t$, i.e., $\mathcal{R}_{t}=f(\mathbf{x}^{*})-\mathcal{F}(x_{t})$, where $\mathcal{F}(x_{t})$ is the cumulative best observation (incumbent best) at iteration $t$, as our performance metric. Fig. 2 shows the convergence plots of GINNBO and the baseline algorithm with $\lambda_{\nabla}=0$ for the three benchmark problems. All regret profiles are based on the mean ±$\pm 1$ standard deviation of $\mathcal{R}_{t}$ over 10 random seeds. Fig. 2 suggests that GINNBO of Algorithm 1 outperforms the baseline BO algorithm for both LCB and LogEI AFs across all three benchmarks. It is evident that GINNBO is agnostic to the choice of AF. In all cases, GINNBO with $\lambda_{\nabla}=1$ in the total loss function reaches much smaller values of the instantaneous regret $\mathcal{R}_{t}$ at earlier iteration indices $t$. When gradient information is available, the BNN surrogate is more likely to locate regions that may contain the global minimum, but can also provide more reliable predictions of the unknown objective (as also seen in Fig. 1), in turn enabling less redundant exploration. The extent to which GINNBO outperforms the baseline algorithm varies with the problem dimensionality $D$. The improved performance is more pronounced in higher dimensions. Even in lower dimensions, it becomes more challenging to learn an accurate BNN surrogate in the limit of sample size. Gradient information allows for effectively scaling the available information per each training sample, $D$-fold; hence, enabling more accurate models to be trained. As can be seen, data scarcity in the case of no derivative observations can make the BNN-based search comparable (or worse) to that of random search.

### v2

@@ 1 Introduction
Bayesian optimization (BO) (Močkus, 1974) has been a powerhouse for optimization of functions with unknown or partially known structure in diverse decision-making applications. BO tackles black-box optimization problems by sequentially solving easier optimization sub-problems. This is done by combining a probabilistic surrogate model, representing the objective as a function of decision inputs, and an acquisition function (AF) that, based on the current belief about the objective, suggests points in the design space to query the objective.
BO is traditionally a zeroth-order optimization method, exclusively relying on objective function observations to refine its belief of the objective function. As for surrogate modeling of the objective, Gaussian process (GP) regression is a popular choice (Williams and Rasmussen, 2006). Recent work has demonstrated the benefits of leveraging derivative information in BO via GPs; fully observed gradients can yield $D+1$ extra observations, where $D$ is the input dimensionality, per system query at a given input. This provides useful local information around the observed points, leading to a more accurate posterior for objective function approximation (Wu et al., 2017b). Gradients have been shown to improve regret bounds in BO (Shekhar and Javidi, 2021), or used for combined global and local search in more complex, potentially multi-objective tasks (Ip et al., 2025, 2024). A standard approach of exploiting gradients (or higher-order derivatives) in BO with GPs is by defining a joint kernel that utilizes both function and gradient observations (Eriksson et al., 2018; Wu et al., 2017a). Alternatively, to mitigate challenges due to noisy function and derivative observations, gradients can be directly modeled using a probabilistic model and utilized to enforce optimality conditions in the AF (Makrygiorgos et al., 2023a, b).
@@ 2.2 Bayesian Neural Networks
$p(\phi(\mathbf{x}),\mathcal{D})=\int p(\phi(\mathbf{x})|\mathbf{x},\theta)p(\theta|\mathcal{D})\,d\theta.$ (4)
@@ 4 Case Studies
We define $\mathcal{R}_{t}$ as the instantaneous regret at iteration $t$, i.e., $\mathcal{R}_{t}=f(\mathbf{x}^{*})-\mathcal{F}(x_{t})$, where $\mathcal{F}(x_{t})$ is the cumulative best observation (incumbent best) at iteration $t$, as our performance metric. Fig. 2 shows the convergence plots of GINNBO and the baseline algorithm with $\lambda_{\nabla}=0$ for the three benchmark problems. All regret profiles are based on the mean $\pm 1$ standard deviation of $\mathcal{R}_{t}$ over 10 random seeds. Fig. 2 suggests that GINNBO of Algorithm 1 outperforms the baseline BO algorithm for both LCB and LogEI AFs across all three benchmarks. It is evident that GINNBO is agnostic to the choice of AF. In all cases, GINNBO with $\lambda_{\nabla}=1$ in the total loss function reaches much smaller values of the instantaneous regret $\mathcal{R}_{t}$ at earlier iteration indices $t$. When gradient information is available, the BNN surrogate is more likely to locate regions that may contain the global minimum, but can also provide more reliable predictions of the unknown objective (as also seen in Fig. 1), in turn enabling less redundant exploration. The extent to which GINNBO outperforms the baseline algorithm varies with the problem dimensionality $D$. The improved performance is more pronounced in higher dimensions. Even in lower dimensions, it becomes more challenging to learn an accurate BNN surrogate in the limit of sample size. Gradient information allows for effectively scaling the available information per each training sample, $D$-fold; hence, enabling more accurate models to be trained. As can be seen, data scarcity in the case of no derivative observations can make the BNN-based search comparable (or worse) to that of random search.